Assume there are 128 routers with 3 lines to other routers.  If costs are represented by 32 bit numbers and distance vectors are exchanged twice per second, how much bandwidth is used by the distance vector routing algorithm?  Assume lines are full duplex.
Answer #1

Answer:

Total number of routers in the network=128

Costs/ delay representation size = 32 bit

The total size of our delay field would be = 128*32 = 4096 bits

As the vectors are exchanged 2 times in a second the field basically updates twice in a second hence we need -

2 * 4096 = 8192 bps (bits per second ) in each direction

As the router is connected to others we will have full utilization hence we have 8192 bps on each line in each direction.